# Dream, Lift, Animate: From Single Images to Animatable Gaussian Avatars

### v1

@@ 3.3 Animate: Structured Gaussian Generation and Deformation
Figure 4: The Gaussian Parameter Decoder (GPD, Sec. 4) maps a UV-aligned latent $\mathbf{Z}$ to an animatable 3D Gaussian representation. The GPD upsamples the avatar latent code $\mathbf{Z}$ and produces two output maps: a canonical Gaussian map $\mathbf{F}_{c}$ and an offset map $\mathbf{F}_{\Delta}$. The offset map $\mathbf{F}_{\Delta}$ adds pose- and view-dependent offsets to the canonical Gaussian $\mathbf{F}_{c}$, enabling pose- and view-dependent effects. Given a pose $\Theta$ and a camera $\pi$, the Gaussians are deformed with linear blend skinning pavlakos2019smplx and rasterized to an RGBA image kerbl20233d .
Gaussian Parameter Decoder. Building upon the UV-aligned latent avatar code $\mathbf{Z}$ produced in the previous stage, we now decode this representation into a fully animatable Gaussian-based avatar. As shown in Fig. 4, the Gaussian Parameter Decoder (GPD) maps the avatar latent code $\mathbf{Z}$ to a UV space Gaussian parameter map, from which structured 3D Gaussians can be sampled. Since convolutional networks greatly benefit from pixel-aligned input features, we design the GPD as a spatially-adaptive ConvNet park2019SPADE ; Zhu_2020_CVPR . The GPD is conditioned on the UV-aligned latent $\mathbf{Z}$ and a one-hot encoded UV segmentation map pavlakos2019smplx . We condition with spatially-adaptive normalization park2019SPADE because pixelwise normalization has been shown to provide great results for generating images that are aligned to semantic maps park2019SPADE ; Zhu_2020_CVPR . At the highest resolution, the GPD branches out into a canonical branch and a pose- and view-dependent branch. The pose- and view-dependent branch receives additional inputs with information about the target geometry (surface normals), viewpoint (Plucker rays jia2020plucker ), and pose (relative vertex position map w.r.t. to the neutral pose). These inputs are created by instantiating the SMPL-X body model with the given target pose and rasterizing its relative vertex location, normals, and plucker rays to the UV space. The inputs are again encoded via spatially adaptive normalization and condition the pose- and view-dependent branch. The two branches are merged with an addition. The output is a UV space Gaussian parameter map $\mathbf{F}\in\mathbb{R}^{H_{G}\times W_{G}\times 14}$, which contains the 3D Gaussian parameters, including alphas, colors, as well as positions, rotations and scales defined in the UV tangent space. The two-branch design enables fast inference by caching the canonical branch. Please refer to Sec. 4.2 for more information.
@@
and rasterize the avatar using Gaussian Splatting kerbl20233d .
@@ 3.4 Training
$\centering\mathcal{L}_{\text{GPD}}=\lambda_{\text{L1}}\mathcal{L}_{\text{L1}}+\lambda_{\text{VGG}}\mathcal{L}_{\text{VGG}}+\lambda_{\text{Mask}}\mathcal{L}_{\text{Mask}}+\lambda_{\text{GAN}}\mathcal{L}_{\text{GAN}}+\lambda_{\text{KL}}\mathcal{L}_{\text{KL}}+\lambda_{\text{C}}\mathcal{L}_{\text{C}},\@add@centering$ (2)
where the L1 reconstruction loss $\mathcal{L}_{\text{L1}}=\lVert\mathbf{I}^{*}-\mathbf{\hat{I}}\rVert_{1}$ measures the absolute difference between the rendered output and ground truth image. For mask supervision, we compute $\mathcal{L}_{\text{Mask}}=\lVert\mathbf{M}^{*}-\mathbf{\hat{A}}\rVert_{1}$, which enforces consistency between the alpha maps from Gaussian splatting $\mathbf{\hat{A}}$ and the ground truth masks $\mathbf{M}^{*}$. To enhance perceptual quality, we incorporate a GAN loss $\mathcal{L}_{\text{GAN}}$ using a Patch Discriminator isola2017image with least squares optimization. The perceptual VGG loss $\mathcal{L}_{\text{VGG}}$ leverages an AlexNet krizhevsky2017imagenet backbone with features masked by the ground truth mask $\mathbf{M}^{*}$. For regularization, we apply a Kullback-Leibler Divergence term $\mathcal{L}_{\text{KL}}$ to constrain the avatar latent maps, and $\mathcal{L}_{\text{C}}=\lVert\mathbf{F}_{\Delta}\rVert_{2}$ encourages minimal offsets in the Gaussian offsets map.
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Input Novel Views and Poses
Figure 5: We animate single input images from SHHQ fu2022styleganhuman and render novel views while changing the body pose, demonstrating the robustness of our method to in-the-wild inputs. See Fig. 1 for more examples.
@@ 4 Experiments
We compare our approach with state-of-the-art methods dreamgaussian ; ho2024sith ; zhang2023sifu ; zhuang2024idolinstantphotorealistic3d for human reconstruction from a single image, demonstrating superior performance in both novel view synthesis and animation tasks (Tbl. 1 and Fig. 6). We then showcase the versatility of our framework through various applications like animation (Fig. 5), editing and interpolation (Fig. 11). Finally, we conduct extensive ablation studies (Tbl. 2) to validate our design choices and analyze the impact of different components on the overall performance.
@@ 4.1 Comparison with State-of-the-art
We compare with several the state-of-the-art for one-shot human reconstruction zhuang2024idolinstantphotorealistic3d ; ho2024sith ; zhang2023sifu ; dreamgaussian . Table 1 presents metrics for novel view synthesis, animation, and their combination. We measure perceptual quality (LPIPS) lpips , structural similarity (SSIM), and photometric accuracy (PSNR). Fig. 7 demonstrates animation capabilities while Fig. 6 shows visual results for novel view synthesis.
Novel view Novel pose Novel view & pose
DreamGaussian dreamgaussian 0.1514 19.48 0.8837 - - - - - -
SiTH ho2024sith 0.1577 18.88 0.8764 - - - - - -
SIFU zhang2023sifu 0.1408 19.42 0.8823 - - - - - -
IDOL zhuang2024idolinstantphotorealistic3d * 0.0696 24.48 0.9261 0.0940 22.80 0.9194 0.1022 22.26 0.9136
Ours 0.0580 25.58 0.9279 0.0471 26.41 0.9351 0.0624 25.09 0.9254
Table 1: Quantitative results on ActorsHQ isik2023humanrf . We compare novel view synthesis on the input image, novel poses using the input camera, and novel views and novel poses. Only IDOL is readily animatable—the other related works would require postprocessing like rigging the reconstructed mesh. Please see Fig. 6 for visuals. The asterik (*) indicates concurrent work.
@@
Figure 6: Comparison for novel view synthesis with DreamGaussian dreamgaussian , SiTH ho2024sith , SIFU zhang2023sifu , and IDOL li2023instant3d . Please see Table 1 for metrics.
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Input IDOL zhuang2024idolinstantphotorealistic3d Ours
Figure 7: Comparison for animation with IDOL zhuang2024idolinstantphotorealistic3d (concurrent SOTA work). Our Dream, Lift, Animate framework enables detailed renderings for difficult poses, outperforming the state-of-the-art in one-shot animatable avatars in perceptual and photometric metrics. Please see Table 1 for metrics.
Our quantitative evaluation in Table 1 demonstrates superior performance in both novel view synthesis and re-posing tasks. For novel view synthesis, we compare our approach against state-of-the-art methods by taking a single image as input and rendering it from multiple unseen viewpoints, as visualized in Fig. 6. The metrics reported in Table 1 confirm our method’s effectiveness across all evaluation criteria. In the following, we discuss the results taking the limitations of the state-of-the-art into consideration.
DreamGaussian dreamgaussian employs score distillation sampling poole2022dreamfusion to produce static Gaussians that can be converted into a mesh with texture refinement, which does not account for view-dependent effects. SiTH ho2024sith and SiFU zhang2023sifu also reconstruct textured meshes using implicit functions, but require additional rigging (e.g., with Mixamo) for animation. IDOL zhuang2024idolinstantphotorealistic3d , a concurrent work, offers fast inference but sacrifices quality, as it directly predicts a Gaussian map in the UV space. This presents two drawbacks: first, it attempts to solve two complex steps (Dream and Lift) simultaneously; second, their Gaussians are not conditioned on the target pose, ignoring pose-dependent effects. These limitations result in reduced reconstruction quality, as evidenced in Table 1 and Fig. 6.
Furthermore, we showcase our animation capabilities on ActorsHQ isik2023humanrf in Fig. 7 and on in-the-wild image fu2022styleganhuman in Figures 1 and 5. The results demonstrate high-quality pose transfer while preserving the avatar’s geometric and appearance. Please see the supplementary material for a comparison on the 4D-Dress dataset wang20244ddress .
@@ 4.2 Applications
Emerging Capabilities. Although our framework is not designed as a generative model, we observe that the learned latent space $\mathbf{Z}$ exhibits smooth structure similar to that of generative style spaces stylegan ; park2019SPADE ; vqvae ; buehler2021varitex ; Rombach2022StableDiffusion . As a result, interpolating between the latent codes of different avatars yields plausible and continuous transitions between identities. Qualitative results are shown in Fig. 8 (top).
@@
Figure 8: Applications. Our structured latent code affords editing (bottom). In addition, we observe emerging capabilities like smooth interpolations between avatar latent codes. These examples are reconstructions using multi-view inputs from CustomHumans ho2023custom .
@@ 4.3 Ablations
We report metrics for several ablation studies in Table 2. The first set of ablations examines techniques for mapping pose-space Gaussians from an unstructured cloud to a structured UV space (A). In ablation A.i, we use the SMPL-X body model to directly project pixel-aligned Gaussians to the UV space, similar to RGB texture unprojection when reconstructing textured meshes from multi-view inputs. This approach lacks learning capabilities and fails to account for shapes beyond the body model (like clothing), resulting in a significant drop in perceptual quality (27% decrease in LPIPS). We also experiment with learning the query (A.ii) instead of using the vertex position map ($\mathbf{X}$ in Fig. 3) and found it yields worse results. We conclude that the vertex position map introduces valuable prior knowledge from the body model, which is missing in the learned query. The second ablation (B) finds a positive impact of view- and pose-conditional inputs in the Gaussian Parameter Decoder (Fig. 4). The third set of ablations (C) considers different inputs to our full model. For inference, our encoder takes the ground-truth frontal input image $\mathbf{I}_{1}$, three synthesized images $\{\mathbf{I}_{2}^{n},...,\mathbf{I}_{V}^{n}\}$, and estimated SMPL-X parameters as input. We study the effect of noisy SMPL-X parameters in the transformer encoder (C.i), which is particularly relevant for in-the-wild settings where fitting the body model might be challenging. Replacing the ground-truth input image $\mathbf{I}_{1}$ with the reconstruction from the diffusion model slightly reduces performance (C.ii). Conversely, we simulate the potential of next-generation video diffusion models by feeding only ground-truth novel views as input, which leads to a substantial performance gain (C.iii), indicating room for further improvement by just replacing the video diffusion model with an improved version, thanks to our our modular design.
Table 2: Ablation studies (Sec. 4.3).
@@ 5 Conclusion
Acknowledgments We thank Shalini De Mello and Jenny Schmalfuss for their valuable inputs, and Jan Kautz for hosting Marcel’s internship.
@@ Appendix B Supplementary Results
Table 3: Quantitative results on 4D-Dress [60]. We compare novel view synthesis on the input image, novel poses using the input camera, and novel views and novel poses. Please see Fig. 9 for visuals. The (*) indicates concurrent work.
@@
IDOL* 0.0904 23.04 0.9226 0.0956 22.54 0.9135 0.0993 21.99 0.9135

### v2

@@ 3.3 Animate: Structured Gaussian Generation and Deformation
Figure 4: Animate.  The Gaussian Parameter Decoder (GPD, Sec. 4) maps a UV-aligned latent $\mathbf{Z}$ to an animatable 3D Gaussian representation. The GPD upsamples the avatar latent code $\mathbf{Z}$ and produces two output maps: a canonical Gaussian map $\mathbf{F}_{c}$ and an offset map $\mathbf{F}_{\Delta}$. The offset map $\mathbf{F}_{\Delta}$ adds pose- and view-dependent offsets to the canonical Gaussian $\mathbf{F}_{c}$, enabling pose- and view-dependent effects. Given a pose $\Theta$ and camera $\pi$, the Gaussians are deformed with linear blend skinning [46] and rasterized to an RGBA image [32].
Gaussian Parameter Decoder. Building upon the UV-aligned latent avatar code $\mathbf{Z}$ produced in the previous stage, we now decode this representation into a fully animatable Gaussian-based avatar. As shown in Fig. 4, the Gaussian Parameter Decoder (GPD) maps the avatar latent code $\mathbf{Z}$ to a UV space Gaussian parameter map, from which structured 3D Gaussians can be sampled. Since convolutional networks greatly benefit from pixel-aligned input features, we design the GPD as a spatially-adaptive ConvNet [45, 79]. The GPD is conditioned on the UV-aligned latent $\mathbf{Z}$ and a one-hot encoded UV segmentation map [46]. We condition with spatially-adaptive normalization [45] because pixelwise normalization has been shown to provide great results for generating images that are aligned to semantic maps [45, 79]. At the highest resolution, the GPD branches out into a canonical branch and a pose- and view-dependent branch. The pose- and view-dependent branch receives additional inputs with information about the target geometry (surface normals), viewpoint (Plucker rays [30]), and pose (relative vertex position map w.r.t. to the neutral pose). These inputs are created by instantiating the SMPL-X body model with the given target pose and rasterizing its relative vertex location, normals, and plucker rays to the UV space. The inputs are again encoded via spatially adaptive normalization and condition the pose- and view-dependent branch. The two branches are merged with an addition. The output is a UV space Gaussian parameter map $\mathbf{F}\in\mathbb{R}^{H_{G}\times W_{G}\times 14}$, which contains the 3D Gaussian parameters, including alphas, colors, as well as positions, rotations and scales defined in the UV tangent space. The two-branch design enables fast inference by caching the canonical branch. Please refer to Sec. 4.2 for more information.
@@
and rasterize the avatar using Gaussian Splatting [32].
@@ 3.4 Training
$\centering\mathcal{L}_{\text{GPD}}\@add@centering$ $=\lambda_{\text{L1}}\mathcal{L}_{\text{L1}}+\lambda_{\text{VGG}}\mathcal{L}_{\text{VGG}}$ (2)
$+\lambda_{\text{Mask}}\mathcal{L}_{\text{Mask}}+\lambda_{\text{GAN}}\mathcal{L}_{\text{GAN}}$ (3)
$+\lambda_{\text{KL}}\mathcal{L}_{\text{KL}}+\lambda_{\text{C}}\mathcal{L}_{\text{C}},$ (4)
where the L1 reconstruction loss $\mathcal{L}_{\text{L1}}=\lVert\mathbf{I}^{*}-\mathbf{\hat{I}}\rVert_{1}$ measures the absolute difference between the rendered output and ground truth image. For mask supervision, we compute $\mathcal{L}_{\text{Mask}}=\lVert\mathbf{M}^{*}-\mathbf{\hat{A}}\rVert_{1}$, which enforces consistency between the alpha maps from Gaussian splatting $\mathbf{\hat{A}}$ and the ground truth masks $\mathbf{M}^{*}$. To enhance perceptual quality, we incorporate a GAN loss $\mathcal{L}_{\text{GAN}}$ using a Patch Discriminator [29] with least squares optimization. The perceptual VGG loss $\mathcal{L}_{\text{VGG}}$ leverages an AlexNet [34] backbone with features masked by the ground truth mask $\mathbf{M}^{*}$. For regularization, we apply a Kullback-Leibler Divergence term $\mathcal{L}_{\text{KL}}$ to constrain the avatar latent maps, and $\mathcal{L}_{\text{C}}=\lVert\mathbf{F}_{\Delta}\rVert_{2}$ encourages minimal offsets in the Gaussian offsets map.
@@ 4 Experiments
We compare our approach with state-of-the-art methods [59, 20, 77, 80] for human reconstruction from a single image, demonstrating superior performance in both novel view synthesis and animation tasks (Tbl. 1 and Fig. 5). We then showcase the versatility of our framework through various applications like animation (Fig. 1), editing and interpolation (Fig. 7). Finally, we conduct extensive ablation studies (Tbl. 2) to validate our design choices and analyze the impact of different components on the overall performance.
@@ 4.1 Comparison with State-of-the-art
We compare with several the state-of-the-art for one-shot human reconstruction [80, 20, 77, 59]. Tbl. 1 presents metrics for novel view synthesis. We measure perceptual quality (LPIPS) [74], structural similarity (SSIM), and photometric accuracy (PSNR). Fig. 6 demonstrates animation capabilities while Fig. 5 shows visual results for novel view synthesis.
Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
DreamGaussian [59] 0.1514 19.48 0.8837
SiTH [20] 0.1577 18.88 0.8764
SIFU [77] 0.1408 19.42 0.8823
IDOL [80] 0.0696 24.48 0.9261
Ours 0.0580 25.58 0.9279
Table 1: Quantitative results on ActorsHQ [28]. We compare novel view synthesis on the input image. Please see Fig. 5 for visuals. The supplementary contains comparisons for novel pose synthesis and comparisons on the 4D-Dress dataset [63].
@@
Figure 5: Comparison for novel view synthesis with DreamGaussian [59], SiTH [20], SIFU [77], and IDOL [35]. Tbl. 1 lists metrics.
@@
Input IDOL [80] Ours
Figure 6: Comparison for animation. Our DLA framework enables detailed renderings for difficult poses, outperforming the state-of-the-art in one-shot animatable avatars in perceptual and photometric metrics. Please see the supp. mat. for more examples and metrics.
Figure 7: Applications. Our structured latent code affords editing like face swapping and virtual try-on of shoes (right). In addition, we observe emerging capabilities like smooth interpolations between avatar latent codes (left). These examples are reconstructions using multi-view inputs from CustomHumans [19].
Our quantitative evaluation in Tbl. 1 demonstrates superior performance in novel view synthesis. We compare our approach against state-of-the-art methods by taking a single image as input and rendering it from multiple unseen viewpoints, as visualized in Fig. 5. The metrics reported in Tbl. 1 confirm our method’s effectiveness across all evaluation criteria. In the following, we discuss the results taking the limitations of the state-of-the-art into consideration.
DreamGaussian [59] employs score distillation sampling [48] to produce static Gaussians that can be converted into a mesh with texture refinement, which does not account for view-dependent effects. SiTH [20] and SiFU [77] also reconstruct textured meshes using implicit functions, but require additional rigging (e.g., with Mixamo) for animation. IDOL [80] offers fast inference but sacrifices quality, as it directly predicts a Gaussian map in the UV space. This presents two drawbacks: first, it attempts to solve two complex steps (Dream and Lift) simultaneously; second, their Gaussians are not conditioned on the target pose, ignoring pose-dependent effects. These limitations result in reduced reconstruction quality, as evidenced in Tbl. 1 and Fig. 5.
Furthermore, we showcase our animation capabilities on ActorsHQ [28] in Fig. 6 and on in-the-wild image [13] in Figures 1. The results demonstrate high-quality pose transfer while preserving the avatar’s geometric and appearance. Please see the supplementary material for a comparison for novel pose synthesis, comparisons on the 4D-Dress dataset [63], and experimental details regarding the comparisons.
@@ 4.2 Applications
Emerging Capabilities. Although our framework is not designed as a generative model, we observe that the learned latent space $\mathbf{Z}$ exhibits smooth structure similar to that of generative style spaces [31, 45, 60, 5, 51]. As a result, interpolating between the latent codes of different avatars yields plausible and continuous transitions between identities. Qualitative results are shown in Fig. 7 (top).
@@ 4.3 Ablations
We report metrics for several ablation studies in Tbl. 2. The first set of ablations examines techniques for mapping pose-space Gaussians from an unstructured cloud to a structured UV space (A). In ablation A.i, we use the SMPL-X body model to directly project pixel-aligned Gaussians to the UV space, similar to RGB texture unprojection when reconstructing textured meshes from multi-view inputs. This approach lacks learning capabilities and fails to account for shapes beyond the body model (like clothing), resulting in a significant drop in perceptual quality (27% decrease in LPIPS). We also experiment with learning the query (A.ii) instead of using the vertex position map ($\mathbf{X}$ in Fig. 3) and found it yields worse results. We conclude that the vertex position map introduces valuable prior knowledge from the body model, which is missing in the learned query. The second ablation (B) finds a positive impact of view- and pose-conditional inputs in the Gaussian Parameter Decoder (Fig. 4). The third set of ablations (C) considers different inputs to our full model. For inference, our encoder takes the ground-truth frontal input image $\mathbf{I}_{1}$, three synthesized images $\{\mathbf{I}_{2}^{n},...,\mathbf{I}_{V}^{n}\}$, and estimated SMPL-X parameters as input. We study the effect of noisy SMPL-X parameters in the transformer encoder (C.i), which is particularly relevant for in-the-wild settings where fitting the body model might be challenging. Replacing the ground-truth input image $\mathbf{I}_{1}$ with the reconstruction from the diffusion model slightly reduces performance (C.ii). Conversely, we simulate the potential of next-generation video diffusion models by feeding only ground-truth novel views as input, which leads to a substantial performance gain (C.iii), indicating room for further improvement by just replacing the video diffusion model with an improved version, thanks to our our modular design. We provide more detailed ablations in the supp. mat.
@@ A.1 Implementation Details
Inference-time Multi-view Generation
As described in the main paper, we synthesize novel views using a ControlNet-guided variant [64] of a video diffusion model [62]. Specifically, we first estimate the SMPL-X parameters from the input image [18]. We then render 2D skeletal poses of the predicted mesh from virtual cameras placed around a 360-degree azimuth with a 0-degree elevation angle. These projected poses serve as control signals to guide the diffusion model in generating photorealistic images from novel viewpoints. We render a video of a 360-degree azimuth rotation in 80 frames and pick the frames 20, 40, 60, and 80 as inputs to our model. For in-the-wild inputs, where major self-contact is common (e.g., hands hidden in the pocket), we condition on an A-pose skeleton. As in the original video diffusion model [64], we use DDIM sampling [56] with 50 steps.
@@ Appendix B Supplementary Results
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Input Novel Views and Poses
Figure 8: We animate single input images from SHHQ [13] and render novel views while changing the body pose, demonstrating the robustness of our method to in-the-wild inputs. See the teaser figure in the main paper for more examples.
Input DreamGaussian SiTH SIFU IDOL Ours GT
Figure 9: Comparison for novel view synthesis with DreamGaussian [59], SiTH [20], SIFU [77], and IDOL [35]. Please see Table 3 for metrics.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Input IDOL [80] Ours
Figure 10: Comparison for animation with IDOL [80]. Our Dream, Lift, Animate framework enables detailed renderings for difficult poses, outperforming the state-of-the-art in one-shot animatable avatars in perceptual and photometric metrics. Please see Table 3 for metrics.
@@
DreamGaussian [59] 0.1514 19.48 0.8837 - - - - - -
SiTH [20] 0.1577 18.88 0.8764 - - - - - -
SIFU [77] 0.1408 19.42 0.8823 - - - - - -
IDOL [80] 0.0696 24.48 0.9261 0.0940 22.80 0.9194 0.1022 22.26 0.9136
Ours 0.0580 25.58 0.9279 0.0471 26.41 0.9351 0.0624 25.09 0.9254
Table 3: Quantitative results on ActorsHQ [28]. We compare extend the comparison from the main paper on novel view synthesis with novel poses using the input camera, and novel views and novel poses jointly. Only IDOL is readily animatable—the other related works would require postprocessing like rigging the reconstructed mesh. Please see Fig. 9 for visuals.
Novel view Novel pose Novel view & pose
IDOL 0.0904 23.04 0.9226 0.0956 22.54 0.9135 0.0993 21.99 0.9135
@@
Table 4: Quantitative results on 4D-Dress [63]. We compare novel view synthesis on the input image, novel poses using the input camera, and novel views and novel poses. Please see Fig. 11 for visuals.
@@
Figure 13: View-dependent effects. The Gaussian Parameter Decoder (GPD) conditions on the view direction with Plucker ray maps (Sec. 3.3 and Fig. 4 in the main paper). This enables view-dependent effects. We render two subjects in an A-pose using the same camera angle but feeding Plucker rays for different view direction: the actual view direction (A), and side and front views (B. and C.). Note the actual view direction (A.) yields sharp details and reflections and other view directions (B. and C.) show effects like a halo on the arms and legs.
Figure 14: Pose-dependent effects. The Gaussian Parameter Decoder (GPD) conditions on the target pose via a relative vertex position map and surface normals (Sec. 3.3 and Fig. 4 in the main paper), enabling pose-dependent effects. We render two subjects in novel poses and visualize the pose-dependent effects, most notable on the shoulders and arms. Please see the website for an animated version.
Figure 15: Applications. Our structured latent code affords editing (bottom). In addition, we observe emerging capabilities like smooth interpolations between avatar latent codes. These examples are reconstructions using multi-view inputs from CustomHumans [19].